OpenStudy (anonymous):

what is the inverse of f(x)=2x-3?

OpenStudy (anonymous):

f(x) = (x - 3)/2

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

it's against the rules to give out answers as they don't help you learn besides, that answer is wrong anyway

OpenStudy (anonymous):

assume y = f(x) = 2x - 3 then calculate x = (y+3)/2 by definition if f(x) = y then y-1=x=f(y) therefore f-1(x) = f(y) = (y+3)/2

OpenStudy (anonymous):

answer is y+3 /2